OpenStudy (thebadyboy):

The total snowfall in Burke, Virginia, this January was 7 1/4 in. In February, 20% more snow fell than in January. How much snow fell in February in Burke, Virginia? Express your answer as a mixed number in simplest form. in.

OpenStudy (anonymous):

To solve this, first we turn 7 1/4 into a number we can get the fraction of; so, 7.25. Next, we times 7.25 by 0.20: \[7.25 * 0.20 = 1.45\] So, we add the 1.45 onto the 7.25: \[7.25 + 1.45 = 8.7\] So, the amount of snow that fell in February was 8 7/10 in. I'm not sure if it can be simplified any further.

OpenStudy (yolo_king):

Nice pro pic xD

OpenStudy (agl202):

First you convert the mixed number (7 1/4) into a normal fraction. Multiply 7 by the denominator, then add the 1 up top, so the fraction is 29/4 Now it's saying 20% (which can also be expressed as .20) MORE snow fell than 29/4. So it would look like 29/4+(.20 times 29/4) Make sense? :)

OpenStudy (agl202):

So remember when multiplying an integer/decimal, it looks like 294×.201=5.84 At this point you wanna add the answer back to the original, so 5.8+294=34.84 And you can't forget to put it back in a mixed number :) 34.84=8710 So 8 7/10 is your answer, I hope that helps!!

OpenStudy (anonymous):

I think @Agl202 just explained it differently.